# Parameter uncertainties for imperfect surrogate models in the low-noise regime

### v3

@@ D.1 Variational optimization
where we use integration by parts. Aside from the POPS constrained variation, this is a standard variational result which in principle allows for variational minimization of $\pi_{E}({\bm{\Theta}})$ at finite $N$ for both linear and non-linear models. An alternative POPS-constrained variational approach for linear regression at finite $N$ is provided in appendix LABEL:app:variational-gaussian.
@@
This result shows the POPS-constrained variation of the PAC-Bayes loss upper bound converges to the gradient of the squared loss evaluated at a POPS-constrained parameter choice, to within a positive multiplicative constant . As a result, in the underparametrized limit the variational optimal choice of of POPS-constrained parameter ${\bm{\Theta}}_{\mathbf{X}}$ will be given by
@@
Appendix E Implementation of the ansatz in Python
The hypercube projector ${\bf V}\in\mathbb{R}^{R\times P}$ and bounds ${\bf B}\in\mathbb{R}^{R\times 2}$ can be produced with the following python code, where A.T is the transpose of A:
1import numpy as np
2def hypercube(X,y,prior,thresh=1e-8):
3 # X.shape = (N,P), y.shape = (N,)
4 # Inverse of A with prior matrix
5 A_i = np.dot(X.T,X) + prior / y.size
6 # Minimum loss solution
7 T_L = np.linalg.solve(A_i,np.dot(X.T,y))
8 # Minimum loss errors
9 e = y - np.dot(X, T_L)
10 # Leverage of each point
11 X_dot_A = np.linalg.solve(A_i,X.T)
12 h = ( X * X_dot_A ).sum(1)
13 # Pointwise fits, shape = (P,N)
14 T = np.multiply(X_dot_A, e/h )
15 # Find hypercube projector
16 nu,V = np.linalg.eigh(T@T.T)
17 nu_sel = nu<nu.max()*thresh
18 V = V[nu_sel,:] # shape = (R,P)
19 # bounds
20 B = np.array([(T@V.T).max(0),(T@V.T).min(0)]).T
21 return V,B
To sample the ensemble we use uniform random numbers
1import numpy as np
2def resample(V,B,n=20,seed=123):
3 np.random.seed(seed)
4 u = np.random.uniform(\
5 min=B[:,0],
6 max=B[:,1],
7 size=(n,V.shape[0]))
8 return u@V
Appendix F Additional plots of test error prediction
F.1 Low dimensional models
@@
We applied the same procedure as for the polynomical example to a model system with $P=6$ independent features to simulate more realistic datasets. As before, the numerical minimization routine could only treat a small range of $\sigma$ values. Figure 4 shows the resulting parameter distributions. At large $\sigma=2$ the minimizer quickly concentrates all ensemble members on the minimum loss solution; at $\sigma=1$ the ensemble distribution is notably sharper than the POPS-ensemble $\pi^{*}_{E}$, whilst for low sigma $\sigma=1/2$ the minimization only has marginal stability, but becomes closer to the POPS-bounding hypercube $\pi^{*}_{\mathcal{H}}$. In addition, a minimizer initialized with the POPS-ensemble did not change appreciably. These albeit simple results give strong evidence that our ansatz is indeed a stable and highly efficient means to find approximate minimizers of ${\mathcal{G}}$ that would otherwise be intractable.
@@
F.2 Medium dimensional models

### v5

@@ D.1 Variational optimization
where we use integration by parts. Aside from the POPS constrained variation, this is a standard variational result which in principle allows for variational minimization of $\pi_{E}({\bm{\Theta}})$ at finite $N$ for both linear and non-linear models.
@@
This result shows the POPS-constrained variation of the PAC-Bayes loss upper bound converges to the gradient of the squared loss evaluated at a POPS-constrained parameter choice, to within a positive multiplicative constant. As a result, in the underparametrized limit, the variational optimal choice of of POPS-constrained parameter ${\bm{\Theta}}_{\mathbf{X}}$ will be given by
@@
Appendix E Implementation of the ansatz
{algorithm}
[h!] Find POPS-Constrained loss minimizers
1:${\bf X}\in\mathbb{R}^{N\times P}$, ${\bf y}\in\mathbb{R}^{N}$, ${\bm{\Sigma}}_{0}\in\mathbb{R}^{P\times P}$ (prior covariance)
2:${\bf C}\leftarrow{\bf X}^{\top}{\bf X}+\frac{{\bm{\Sigma}}_{0}}{N}$triangleright Compute regularized feature covariance matrix
3:${\bf w}\leftarrow\text{solve}({\bf C},{\bf X}^{\top}{\bf y})$triangleright Solve ${\bf C}{\bf w}={\bf X}^{\top}{\bf y}$ for least squares solution
4:${\bf e}\leftarrow{\bf y}-{\bf Xw}$triangleright Calculate residuals
5:${\bf A}\leftarrow\text{solve}({\bf C},{\bf X}^{\top})$triangleright Solve ${\bf C}{\bf A}={\bf X}^{\top}$ for influence matrix
6:${\bf h}\leftarrow\text{diag}({\bf X}{\bf A})$triangleright Compute vector of leverage scores
7:${\bf T}\leftarrow{\bf A}\odot({\bf e}\oslash{\bf h})$triangleright Element-wise product and division for POPS-constrained corrections to ${\bf w}$
8:return ${\bf w},{\bf T}$triangleright Return the loss minimizer and POPS-constrained corrections
{algorithm}
[h!] Generate POPS-constrained hypercube samples
1:${\bf T}\in\mathbb{R}^{N\times P}$ (POPS-constrained corrections from Algorithm E)
2:${\bf U},{\bf S},{\bf V}^{\top}\leftarrow\text{SVD}({\bf T})$triangleright Perform SVD on T
3:$R\leftarrow\text{rank}({\bf S})$triangleright Determine rank of S
[…]
5:$\tilde{\bf T}\leftarrow{\bf T}{\bf V}_{R}$triangleright Project T onto the singular vectors
6:${\bf l}\leftarrow\min(\tilde{\bf T},\text{axis}=0)$triangleright Find min bounds of projections
7:${\bf u}\leftarrow\max(\tilde{\bf T},\text{axis}=0)$triangleright Find max bounds of projections
8:for $i=1$ to $M$ do $\triangleright$ Generate M samples
9:     ${\bf x}\leftarrow\text{UniformRandom}({\bf l},{\bf u})$triangleright Sample uniformly within bounds
10:     ${\bf T}_{\text{sample},i}\leftarrow{\bf x}{\bf V}_{R}^{\top}$triangleright Project back to original basis
11:end for
12:return ${\bf T}_{\text{sample}}$triangleright Return M hypercube samples
A slightly modified implementation of this POPS-constrained linear regression algorithm, wrapping the Scikit-learn BayesianRidge classPedregosa et al. (2011), is available on GitHub at https://github.com/tomswinburne/POPS-Regression.git. It can be installed (requires Scikit-learn) via pip: pip install POPSRegression.
Figure 6: Comparision of POPSRegression against BayesianRidge from sklearn. Left: Mean and standard deviation of the time taken for each method to fit a $P$ dimensional linear model against $N=10000$ training points, with $P$ ranging from 10 to 1000. Right: Mean and standard deviation of the time taken to make $1000$ predictions with mean and epistemic uncertainties for BayesianRidge, whilst POPSRegression returns mean, misspecification uncertainty and max/min bounds.
Appendix F Performance comparison against Bayesian Ridge
To demonstrate the performance of POPSRegression against BayesianRidge from sklearn, we fit against the medium-dimensional cubic ”simulation engine” described in the main text, with $P$ ranging from 10 to 1000 and $N=10000$. Figure 6 shows POPSRegression incurs a minimal (2×2$\times$2 ×) overhead over BayesianRidge, whilst existing misspecification aware schemes do not converge in the low-noise regime.
Appendix G Additional plots of test error prediction
G.1 Low dimensional models
@@
We applied the same procedure as for the polynomical example to a model system with $P=6$ independent features to simulate more realistic datasets. As before, the numerical minimization routine could only treat a small range of $\sigma$ values. Figure 7 shows the resulting parameter distributions. At large $\sigma=2$, the minimizer quickly concentrates all ensemble members on the minimum loss solution; at $\sigma=1$ the ensemble distribution is notably sharper than the POPS-ensemble $\pi^{*}_{E}$, whilst for low sigma $\sigma=1/2$ the minimization only has marginal stability, but becomes closer to the POPS-bounding hypercube $\pi^{*}_{\mathcal{H}}$. In addition, a minimizer initialized with the POPS-ensemble did not change appreciably. These albeit simple results give strong evidence that our ansatz is indeed a stable and highly efficient means to find approximate minimizers of ${\mathcal{G}}$ that would otherwise be intractable.
@@
G.2 Medium dimensional models